# date112023

### v1

@@ 1. Introduction
The term “beta”, or $\beta$, represents the linear coefficient between an explanatory variable and a response variable, and plays a crucial role in finance markets for asset pricing, portfolio optimization, and risk management, among other functional areas. In asset pricing, $\beta$ is used to evaluate the return of an asset based on, for example, the return of the market portfolio (Capital Asset Pricing Model, CAPM (Sharpe, 1964)), market plus asset characteristics (Fama-French three-factor model and five-factor model (Fama and French, 1993, 2015), Carhart four-factor model (Carhart, 1997)), or a set of risk factors and market indices (Arbitrage Pricing Theory, APT (Ross, 2013)).
Due to the influence of various market dynamics, $\beta$ evolves over time, posting challenges for estimation. Many methods have been devised based on panel regression, rolling regression, or other time-varying inference. Examples include (Fama and MacBeth, 1973; Groenewold and Fraser, 1999; Faff et al., 2000; Adrian et al., 2015; Jagannathan and Wang, 1996; Lettau and Ludvigson, 2001). Recent studies have also introduced machine learning based estimators such as those by (Jourovski et al., 2020; Drobetz et al., 2023; Gu et al., 2020; Mirete-Ferrer et al., 2022). In general, $\beta$ estimators aim to capture the latest market dynamics by identifying certain patterns in historical data.
In this paper, we propose a novel approach to beta estimation using deep neural networks, which we call NeuralBeta. NeuralBeta leverages the power of neural networks to capture complex, non-linear relationships in financial data, offering more precise and robust estimation of $\beta$. We also develop an interpretable variant, NeuralBeta-Interpretable (NBI), since model interpretability is important for risk management and governance purposes. NBI enhances transparency by explicitly outputting weights for each data point in the lookback window, which are then used to calculate the weighted least squares (WLS) solution to estimate $\beta$. This approach allows users to easily discern which historical periods are more influential in the beta estimation process and to identify specific temporal patterns and market conditions that influence beta.

### v2

@@ 1. Introduction
The term “beta”, or $\beta$, represents the linear coefficient between an explanatory variable and a response variable, and plays a crucial role in finance markets for asset pricing, portfolio optimization, and risk management. In asset pricing, $\beta$ is used to evaluate the return of an asset based on, for example, the return of the market portfolio (Capital Asset Pricing Model (Sharpe, 1964)), market plus asset characteristics (Fama-French three-factor model and five-factor model (Fama and French, 1993, 2015)), or a set of risk factors and market indices (Arbitrage Pricing Theory (Ross, 2013)).
The original CAPM model assumes that stock market betas remain constant over time. However, empirical research including (Bollerslev et al., 1988) and (Jagannathan and Wang, 1996) demonstrates that betas are indeed time-varying. Traditional methods to model this time variation usually involve rolling-window regressions using Ordinary Least Squares (OLS) and Weighted Least Squares (WLS). But these approaches face challenges in specifying the window length and weights to balance bias and variance. These regression methods are also sensitive to outliers, which leads to volatile beta estimates.
In recent years, with the proliferation of machine learning in finance, research applying machine learning techniques to beta estimation remains sparse. (Drobetz et al., 2023), for example, utilizes future information to compute the “true beta” (or realized beta) as the target and evaluates several machine-learning based estimators against this metric, and they found random forests to be the most effective. A limitation of this approach is the introduction of new hyperparameters in computing realized beta (the window length, the weights, etc.) and that using future data in computing the target reduces the amount of data available. Moreover, (Alanis et al., 2024) explores the use of machine learning algorithms to estimate equity betas from historical returns and company financials, for which they also used the realized beta as the target. They focus primarily on private firms or non-traded assets, which diverges from our area of interest. Notably, none of these approaches employ state-of-the-art deep neural networks like attention, nor do they provide the transparency needed in financial modelling. The usage of realized beta as the target is also questionable, as the primary usage of beta is hedging, and targeting a beta using a future window does not necessarily improve this task.
Our paper contributes to the literature by proposing an interpretable approach to beta estimation using deep neural networks, which we call NeuralBeta. NeuralBeta directly uses the hedging error as the metric, and leverages the power of neural networks to capture complex, non-linear relationships in financial data, offering more precise and robust estimation of $\beta$. We also develop an interpretable variant, NeuralBeta-Interpretable (NBI). NBI enhances transparency by explicitly outputting weights for each data point in the lookback window, which are then used to calculate the weighted least squares (WLS) solution to estimate $\beta$. This approach allows users to easily discern which historical periods are more influential in the beta estimation process and to identify specific temporal patterns and market conditions that influence beta.
@@ 2.2.2. Interpretable Neural Network Architecture
Naturally, there remains a certain degree of non-interpretability further upstream, such as understanding the factors that lead certain observations to carry more weights than others. However, this approach takes a meaningful step towards interpretability and generically aid in viewing neural networks as estimators; we leave it to future work to study if the interpretability can be further enhanced.